# If quasars form from primordial black holes

### v1

@@ 3.3 Radio galaxies
In hindsight, early results on radio galaxies (Ekers et al. 1981), two-thirds of the sources being double, the merger, Cen A, and VLBI jets pointed to the presence of black holes and the importance of accretion. The variety of radio galaxies called at the same early time for unification (Readhead 2015), whereby diverse properties could be related to the perspective of the observer relative to the axis of the galaxy. In their review Urry & Padovani (1995) asked ten questions to be answered to carry unification further, of which we note five that still seem outstanding and demanding.
Is there evidence for obscured quasars in all radio galaxies?
How do jets form and propagate?
The physical cause of the radio loud - radio quiet distinction?
Where are the narrow-line (Type 2) radio-quiet quasars?
What are the fundamental parameters governing the central engine, and is it powered by a black hole?
Radio galaxies have a similar LF, but scaled down in maximum radio power by 6.35 dex (Pracy et al. 2016). Applying a bolometric correction of 4.5 dex (Marconi et al. 2004 and Panessa et al. 2007), we find correspondence to PBH nuclei of 106.5, overlapping in nuclear mass with QSOs, and implying a qualitative difference in their accretion state. At L46 the QSO LF is dlog n / dlog L = –4.7 = –dn/n t/dt, since log L = 2 log c +log m –log t, and so dn/n t/dt = 4.7. Continuity requires that dn/dt is constant; so, lifetime in state x nx where x = quasar, Seyfert, or radio galaxy, if they are all connected by the same galaxy interaction rate. If QSO and radio galaxy are different stages of the same phenomenon, the lifetime as radio galaxy is 10 times that as quasar, according to the LFs.
We have not fully answered any of the questions noted earlier. But the notion of a pause between accretion episodes, should be added to the unification premise of observer perspective. Lifetimes of different phases and what can be learned from jets are reviewed by Garofolo et al. 2018). Duty cycles of radio galaxies are discussed by Shabala et al. (2020).
@@ 4 Conclusions
Although the caveat was noted by the first proponents of PBH as QSO progenitors, the possibility that formation of 107 M⊙ PBH might leave a mark on Big Bang Nucleosynthesis has not been considered here. Given readily available light element abundance data, plus the large infall radii of SMBH we outlined in 2, the consequences should be explored in detail with calculations of relativistic infall and the resulting density enhancements. These infall regions are galaxy sized. A dispersion in galaxy initial helium abundances, for example, might be expected.
@@
Whether SMBH can pass undetected through the surface of last scattering of the CMB is an open question that could falsify the model via impending sensitive observations.
@@ Acknowledgements
We would like to thank our colleagues in the Swinburne CAS & ANU RSAA microlensing team and the ARC Centre of Excellence for Dark Matter Particle Physics for stimulating discussions, and we acknowledge ARC grant CE200100008 which, together with the five Australian university nodes, funds the centre’s research. We also acknowledge the generous intellectual support of the centre’s international partners. We thank Matthew Bailes for reminding us about the IGM and Elaine Sadler for presenting the radio galaxy luminosity function at the recent centenary of Mt. Stromlo Observatory and Karl Glazebrook for an update on little red dots.
Jeremy Mould adds his thanks to Peter Goldreich for lessons in astrophysical timescales, Phil Hopkins for his calculations and movies of black holes and galaxies interacting and Alvio Renzini who developed the fuel burning theorem in stellar evolution. In memory of our late colleagues at Caltech, Jesse Greenstein, Wal Sargent and Martin Schmidt, quasar pundits all.
Appendix
Supplementary information on equations 1-6
1.
Number density = 1/volume; volume =4/3, where r is the Schwarzschild radius.
2.
Nuclear reaction rate equation, Clayton (1968) eq 6.14.
3.
Energy density = collision rate energy produced per collision.
6.
This luminosity function is obtained by integrating the M7 number density by the m-1 IMF.
In addition to the constraint on M7 number density in equation (5), there is a number density at the time of M7 origin. This derives from the volume occupied by the M7 Schwarzschild radius at formation time, 400 s after the Big Bang of
At a redshift of 7.47 1010 this density becomes

### v2

@@ 3.4 Radio galaxies
In hindsight, early results on radio galaxies (Ekers et al. 1981), two-thirds of the sources being double, the merger, Cen A, and VLBI jets pointed to the presence of black holes and the importance of accretion. The variety of radio galaxies called at the same early time for unification (Readhead 2015), whereby diverse properties could be related to the perspective of the observer relative to the axis of the galaxy.
Radio galaxies have a similar LF, but scaled down in maximum radio power by 6.35 dex (Pracy et al. 2016). Applying a bolometric correction of 4.5 dex (Marconi et al. 2004 and Panessa et al. 2007), we find correspondence to PBH nuclei of 106.5, overlapping in nuclear mass with QSOs, and implying a qualitative difference in their accretion state. At L46 the QSO LF is dlog n / dlog L = –4.7 = –dn/n t/dt, since log L = 2 log c +log m –log t, and so dn/n t/dt = 4.7. Continuity requires that dn/dt is constant; so, lifetime in state x nx where x = quasar, Seyfert, or radio galaxy, if they are all connected by the same galaxy interaction rate. If QSO and radio galaxy are different stages of the same phenomenon, the lifetime as radio galaxy is 10 times that as quasar, according to the LFs.The notion of a pause between accretion episodes, should be added to the unification premise of observer perspective. Lifetimes of different phases and what can be learned from jets are reviewed by Garofolo et al. 2018). Duty cycles of radio galaxies are discussed by Shabala et al. (2020).
@@ 4 Conclusions
Although the caveat was noted by the first proponents of PBH as QSO progenitors, the possibility that formation of 107 M⊙ PBH might leave a mark on Big Bang Nucleosynthesis has not been considered here ( Sanderbeck, Bird, & Haiman, 2021). Given readily available light element abundance data, plus the large infall radii of SMBH we outlined in 2, the consequences should be explored in detail with calculations of relativistic infall and the resulting density enhancements. These infall regions are galaxy sized. A dispersion in galaxy initial helium abundances, for example, might be expected.
@@ Acknowledgements
We would like to thank our colleagues in the Swinburne CAS & ANU RSAA microlensing team and the ARC Centre of Excellence for Dark Matter Particle Physics for stimulating discussions, and we acknowledge ARC grant CE200100008 which, together with the five Australian university nodes, funds the centre’s research. We also acknowledge the generous intellectual support of the centre’s international partners. We thank Matthew Bailes for reminding us about the IGM and Elaine Sadler for presenting the radio galaxy luminosity function at the recent centenary of Mt. Stromlo Observatory and Karl Glazebrook for an update on little red dots. We are grateful to the referees for improving the paper.
Jeremy Mould adds his thanks to Alvio Renzini who developed the fuel burning theorem in stellar evolution. In memory of our late colleagues at Caltech, Jesse Greenstein, Wal Sargent and Martin Schmidt, quasar pundits all.